Find the value of x: logx 121/361=2
Answer by stanbon(48535)

Find the value of x: logx 121/361=2
--------------
Write in exponential form:
x^2 = (121/361)
Take the square root of both sides
x = 11/19 or x = -11/19
===============
Cheers,
Stan H.